### v2

# Wormhole geometries in Einstein-aether theory

Abstract
We present the first analysis of traversable wormhole solutions within the framework of Einstein–aether theory. We show that the corresponding field equations admit three distinct wormhole geometries, obtained by adopting three different classes of combinations for the aether coupling constants. We examine the null and weak energy conditions for three types of wormhole shape functions. Our findings reveal that, in contrast to Einstein gravity, by choosing appropriate parameter values, wormhole geometries can satisfy the energy conditions at the wormhole throat. We also find that in one class, wormholes can satisfy the energy conditions not only at the wormhole throat but also throughout the entire spacetime. Furthermore, the requirement of energy condition satisfaction, imposes some constraints on the values of aether coupling constants. By comparing these constraints with those previously obtained from theoretical and observational analyses, we find that the satisfaction of energy conditions put more stringent limits on the allowed values of the aether couplings.
1 Introduction
A wormhole is a hypothetical topological feature of spacetime that creates shortcuts between two distinct spacetimes or between two distant points within a single spacetime. The first solution describing such a structure was found by Flamm in 1916 in the context of General Relativity (GR) [1]. In 1935, Einstein and Rosen constructed an unstable, non-traversable wormhole, known as the Einstein-Rosen bridge [2]. They discovered a tunnel-like geometry through which a black hole could form a bridge to a remote region of spacetime. The term “wormhole” was first introduced in 1957 in the seminal papers of Misner and Wheeler [3, 4] to describe a mechanism for “charge without charge.” They analyzed the Riemannian geometry of manifolds with nontrivial topology.
The study of Lorentzian wormholes in the context of GR dates back to the foundational work of Morris and Thorne in 1988. Motivated by the possibility of rapid interstellar travel [5, 6], and by introducing a static, spherically symmetric line element, they showed that exact solutions representing wormhole geometries could be found by solving the Einstein field equations. It was also found that these traversable wormholes possess a stress-energy tensor that violates the Null Energy Condition (NEC) and thus, traversable wormholes are only possible if so-called “exotic matter” exists at their throat, involving an energy-momentum tensor that violates the NEC (see, e.g., [7] or [8] for more recent reviews). From a theoretical standpoint, quantum field theory allows for the existence of exotic energy, as demonstrated by the Casimir effect [9]. Wormhole solutions supported by Casimir energy have been extensively studied in Ref. [10]. Exotic matter also appears in cosmological scenarios. For instance, phantom energy exhibits the exotic property of supporting wormhole geometries [11, 12, 13, 14]. For static wormholes in Einstein gravity, the null energy condition is violated; therefore, several attempts have been made to overcome this issue.
One of the most significant challenges in wormhole physics concerns the fulfillment of the standard energy conditions. Consequently, many studies have sought realistic matter sources that can support wormhole configurations or minimize the need for exotic matter. Research in this area has focused on the construction of thin-shell, dynamical, and rotating wormholes [15, 16, 17, 18, 19, 20, 21]. However, in this work, we focus on modified gravity theories, in which the effective framework allows for static, spherically symmetric, and traversable wormhole solutions while reducing their dependence on exotic matter. Higher-order curvature terms have also been shown to allow the construction of thin-shell wormholes supported by ordinary matter [22, 23, 24]. Similar studies have explored wormhole geometries in Brans-Dicke theory [25, 26], $f(R)$ gravity [27, 28, 29], Kaluza-Klein gravity [30, 31], Rastall gravity [32], scalar-tensor gravity [33, 34, 35], solutions in the presence of a cosmological constant [36, 37], non-commutative geometry [38, 39] and other modified gravity theories [40, 41, 32, 42].
In the present work, we aim to consider the Einstein-Aether (EA) theory and investigate the effects of it on the possibility of having wormhole solutions with normal matter. In the following, we briefly review the EA theory and its black hole solutions.
The Einstein–Aether theory (commonly referred to as æ-theory) is a generally covariant modification of general relativity that introduces a dynamical, unit timelike vector field, the “æther”, which establishes a preferred reference frame and thus violates Lorentz boost invariance. This theory was first clearly formulated by Ted Jacobson and collaborators to explore Lorentz-violating effects in the gravitational sector while preserving diffeomorphism invariance [43, 44].
Research shows the existence of static, spherically symmetric black hole solutions in EA theory. These solutions closely resemble the Schwarzschild geometry outside the metric horizon, with deviations typically at the percent level for parameter regions consistent with empirical bounds. Notably, these black holes possess a distinctive universal horizon, a surface deeper than the metric horizon that traps even modes propagating arbitrarily fast relative to the æther [45, 46].
Early studies by Eling and Jacobson revealed that regular, asymptotically flat black hole families exist, provided regularity at both the metric and spin-0 horizons; numerical solutions show oscillatory æther behavior near singularities [47]. Further refinements considering parameterized post-Newtonian constraints demonstrate scenarios where spin-0 horizons lie inside the metric horizon, allowing significant deviations in ADM mass, innermost stable circular orbit, and Hawking temperature-though such features may remain observationally elusive [48]. Dynamic scenarios, including gravitational collapse, have also been numerically simulated by Garfinkle, Eling, and Jacobson, confirming that black holes form with regular æther configurations under typical conditions [49].
The ringdown properties of black holes in Einstein–æther theory have been investigated, revealing that quasinormal modes are modified: damping rates generally decrease and ringing periods increase relative to GR, offering possible, but subtle, opportunities for gravitational-wave probes [50, 51]. Studies of quantum tunneling at both Killing and universal horizons in Einstein-Maxwell-æther black holes indicate nontrivial effects on Hawking radiation, especially under higher-order curvature corrections and modified dispersion relations [52].
The organization of this paper is as follows. In Sec. 2, we present a brief review of the EA theory and its corresponding field equations. In Sec. 3, we investigate traversable wormhole geometries within the framework of EA theory. By adopting an anisotropic form for the energy–momentum tensor, we reformulate the field equations accordingly. In Secs. 4, 5, and 6, we obtain solutions to these field equations for three different classes of combinations of the EA coupling constants. Within each class, we analyze three types of wormhole shape functions, and after determining the energy density and pressure components, we examine the energy conditions associated with the resulting wormhole solutions. Finally, in Sec. 7, we summarize and discuss our main findings.
2 Einstein-Aether theory and Field equations
The action of EA theory is defined as [43, 53]
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{16\pi G}\left(R+\mathcal{L}_{\ae }\right)+\mathcal{L}_{M}\right],$ (2.1)
where
$\mathcal{L}_{\ae }=-{K^{ab}}_{mn}\nabla_{a}u^{m}\nabla_{b}u^{n}+\lambda(g_{ab}u^{a}u^{b}+1),$ (2.2)
and
${K^{ab}}_{mn}=c_{1}g^{ab}g_{mn}+c_{2}\delta^{a}_{m}\delta^{b}_{n}+c_{3}\delta^{a}_{n}\delta^{b}_{m}-c_{4}u^{a}u^{b}g_{mn}.$ (2.3)
Here $c_{i}$’s are dimensionless aether coupling constants, $u^{a}$ is the aether field, $\lambda$ is a Lagrange multiplier enforcing the unit time-like constraint on the aether, and $\mathcal{L}_{M}$ is the matter Lagrangian which depends on the metric and matter fields. The Kronecker delta is defined as
$\delta^{a}_{m}=g^{a\alpha}g_{\alpha m}\,.$ (2.4)
Variation of the action (2.1) with respect to $\lambda$ yields the constraint
$g_{ab}u^{a}u^{b}=-1\,$, (2.5)
which ensures that the aether field is a unit time-like vector. In the weak-field and slow-motion limit, the EA theory reduces to Newtonian gravity, where the Newtonian constant $G_{N}$ is related to $G$ in (2.1) as [49]
$G=G_{N}\big(1-\frac{c_{14}}{2}\big)\,$, (2.6)
with $c_{ij}\equiv c_{i}+c_{j}$. Observational and theoretical considerations impose the following constraints on the coupling constants [54, 55, 53]
$0\leq c_{14}<2\,,\qquad 2+c_{13}+3c_{2}>0\,,\qquad 0\leq c_{13}<1\,.$ (2.7)
Variation of the action (2.1) with respect to the aether field $u^{a}$ gives the aether field equation [49]
$\nabla_{a}J^{a}_{b}+c_{4}a_{a}\nabla_{b}u^{a}+\lambda u_{b}=0,$ (2.8)
where
$J^{a}_{~m}={K^{ab}}_{mn}\nabla_{b}u^{n},\qquad a_{a}=u^{b}\nabla_{b}u_{a}\,$, (2.9)
Variation of (2.1) with respect to the metric $g_{\mu\nu}$ leads to [56, 53, 57]
$\mathcal{G}_{ab}=T_{ab}^{\ae }+8\pi GT^{M}_{ab}\,$, (2.10)
where $\mathcal{G}_{ab}=R_{ab}-\frac{1}{2}Rg_{ab}$ is the Einstein tensor, and the aether energy-momentum tensor is
$T_{ab}^{\ae }=\lambda u_{a}u_{b}+c_{4}a_{a}a_{b}-\frac{1}{2}g_{ab}J^{c}_{~d}\nabla_{c}u^{d}+\nabla_{c}X^{c}_{~ab}+c_{1}[(\nabla_{a}u_{c})(\nabla_{b}u^{c})-(\nabla^{c}u_{a})(\nabla_{c}u_{b})],$
$~~~~~~~~~~~~X^{c}_{~ab}=J^{c}_{~~(a}u_{b)}-u_{(a}J^{~~c}_{b)}+u^{c}J_{(ab)}\,,\qquad\lambda=c_{4}a^{2}+u^{a}\nabla_{b}J^{b}_{~a}\,.$ (2.11)
The matter energy-momentum tensor is also defined as
$T^{M}_{ab}=-\frac{2}{\sqrt{-g}}\frac{\delta(\sqrt{-g}\mathcal{L}_{M})}{\delta g_{ab}}.$ (2.12)
For a static and spherically symmetric spacetime
$ds^{2}=-A(r)dt^{2}+B(r)dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2},$ (2.13)
the field equations can be solved for certain choices of the coupling constants $c_{i}$’s [56]. In particular, the following classes of solutions have been identified for black holes. I: $c_{2}\neq 0$, $c_{13}\neq 0$, and $c_{14}=0$, II: $c_{2}\neq 0$, $c_{13}=0$, and $c_{14}=0$, III: $c_{2}=0$, $c_{13}\neq 0$, and $c_{14}=0$, IV: $c_{2}=0$, $c_{13}=0$, and $c_{14}\neq 0$, $\cdots$. In the next section, we focus on traversable wormhole solutions in EA theory and analyze their energy conditions.
3 Treversable wormhole solutions
Let us start with the line element of a traversable wormhole as in [5]
$ds^{2}=-dt^{2}+\frac{dr^{2}}{1-\frac{b(r)}{r}}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)\,$, (3.1)
where $b(r)$ is the wormhole shape function, which satisfies the following conditions at the wormhole throat $r_{0}$
$b(r_{0})=r_{0}\,$, (3.2)
$b^{\prime}(r_{0})<1\,.$ (3.3)
To preserve the metric signature throughout the wormhole spacetime $r>r_{0}$, it is also necessary that
$1-\frac{b(r)}{r}>0\,.$ (3.4)
To study wormhole solutions in the background of EA theory, we need to specify the aether field $u^{a}$. Due to the static and spherically symmetric structure of the metric (3.1), we adopt the following aether field ansatz [56]
$u^{a}=\left[a(r),h(r),0,0\right]\,.$ (3.5)
Noting the constraint (2.5) it is straightforward to express $h(r)$ in terms of $a(r)$ as
$h(r)=\sqrt{\frac{\left[r-b(r)\right]\left[a(r)^{2}-1\right]}{r}}\,.$ (3.6)
By plugging the metric (3.1) and aether field (3.5) into the field equations (2.8)-(2.12) and considering a diagonal energy-momentum tensor for the matter field as ${T^{a}_{~b}}^{M}={\rm diag}[-\rho(r),p_{r}(r),p_{t}(r),p_{t}(r)]$, where $\rho$ is the energy density and $p_{r}$, $p_{t}$ are the radial and transverse pressure, respectively, one obtains the following equations
$\frac{c_{2}}{2r^{3}(a^{2}\!-\!1)}\bigg\{2ar^{2}(a^{2}\!-\!1)^{2}(r\!-\!b)a^{\prime\prime}-3(r\!-\!b)r^{2}(a^{2}\!-\!\frac{2}{3})a^{\prime 2}-ra(a^{2}\!-\!1)a^{\prime}[r(a^{2}\!-\!1)b^{\prime}$
$+(3b\!-\!4r)a^{2}+8r-7b]-2(a^{2}\!-\!1)^{2}\left[r(a^{2}\!-\!1)b^{\prime}+a^{2}(2r-3b)+b\right]\bigg\}$
$+\frac{c_{13}}{2r^{3}(a^{2}\!-\!1)}\bigg[2ar^{2}(a^{2}\!-\!1)(r\!-\!b)a^{\prime\prime}\!-3(r\!-\!b)r^{2}(a^{2}\!-\!\frac{2}{3})a^{\prime 2}-ra(a^{2}\!-\!1)^{2}(b^{\prime}r\!-\!4r\!+3b)a^{\prime}$
$-4(a^{2}\!-\!1)^{2}(r\!-\!b)(a^{2}\!-\!\frac{1}{2})\bigg]-\frac{c_{14}}{2r^{2}}\bigg\{2ar(a^{2}\!-\!2)(r\!-\!b)a^{\prime\prime}-\big[r(r\!-\!b)a^{\prime}$
$+a(a^{2}\!-\!2)(b^{\prime}r\!-\!4r\!+3b)\big]a^{\prime}\bigg\}-\frac{b^{\prime}}{r^{2}}+\rho=0\,$, (3.7)
$\frac{c_{2}}{2r^{3}(a^{2}\!-\!1)}\bigg\{-2a^{3}r^{2}(a^{2}\!-\!1)(r\!-\!b)a^{\prime\prime}+a^{2}r^{2}(r\!-\!b)a^{\prime 2}+(a^{2}\!-\!1)[ra^{2}b^{\prime}+(3b\!-\!4r)a^{2}$
$-4r+4b]raa^{\prime}+2(a^{2}\!-\!1)^{2}\left[ra^{2}b^{\prime}+a^{2}(2r-3b)-2r+2b\right]\bigg\}$
$+\frac{c_{13}}{2r^{3}(a^{2}\!-\!1)}\bigg[-2a^{3}r^{2}(a^{2}\!-\!1)(r\!-\!b)a^{\prime\prime}+a^{2}r^{2}(r\!-\!b)a^{\prime 2}+(a^{2}\!-\!1)(3b\!-\!4r\!+\!b^{\prime}r)ra^{3}a^{\prime}$
$+4(a^{2}\!-\!1)^{2}(r\!-\!b)(a^{2}\!-\!\frac{1}{2})\bigg]+\frac{c_{14}}{2r^{2}}\bigg\{2ar(a^{2}\!-\!1)(r\!-\!b)a^{\prime\prime}-\big[r(r\!-\!b)a^{\prime}$
$-a(a^{2}\!-\!1)(b^{\prime}r\!-\!4r\!+3b)\big]a^{\prime}\bigg\}-\frac{b}{r^{3}}-p_{r}=0\,$, (3.8)
$\frac{c_{2}}{2r^{3}(a^{2}\!-\!1)}\big[-2ar^{2}(a^{2}\!-\!1)(r\!-\!b)a^{\prime\prime}-r^{2}(a^{2}\!-\!2)(r\!-\!b)a^{\prime 2}+ar(a^{2}\!-\!1)(b^{\prime}r\!-\!8r+\!7b)a^{\prime}$
$+2(a^{2}\!-\!1)^{2}(b^{\prime}r-b)\big]+\frac{c_{13}}{2r^{3}(a^{2}\!-\!1)}\big[a^{2}r^{2}(r\!-\!b)a^{\prime 2}-4ra(a^{2}\!-\!1)(r\!-\!b)a^{\prime}$
$+(a^{2}\!-\!1)^{2}(b^{\prime}r\!-\!b)\big]-\frac{c_{14}}{2r}(r-b)a^{\prime 2}+\frac{b-b^{\prime}r}{2r^{3}}-p_{t}=0\,.$ (3.9)
Here, a prime denotes differentiation with respect to $r$. Our goal is to investigate the energy conditions for wormhole solutions in this theory. To this end, we first obtain explicit expressions for the energy density $\rho$ and pressure components $p_{r}$ and $p_{t}$, and then examine the validity of the Weak Energy Condition (WEC) by simultaneously checking the following inequalities: EC1: $\rho>0$,  EC2: $\rho+p_{r}>0$ and EC3: $\rho+p_{t}>0$. We also verify the NEC by examining EC2 and EC3. Note that if the WEC holds, the NEC is automatically satisfied.
The equations (3)-(3) are solvable for certain values of the coupling constants $c_{i}$. Among these, we identify three classes of solutions in which the satisfaction of the energy conditions imposes constraints on the aether coupling constants. In other words, we focus on traversable wormhole solutions in the EA theory that satisfy the energy conditions. Our calculations show that such solutions exist if the coupling constants lie within specific intervals. In the following sections, we study these classes of wormhole solutions by considering three types of shape functions $b(r)$ and investigate the corresponding energy conditions.
4 Class I: $c_{13}=0$, $c_{14}=0$, and $c_{2}\neq 0$
By choosing the coupling constants of the EA theory in this manner, the equations (3)-(3) become solvable, yielding [56]
$a(r)=\frac{d}{r^{2}}.$ (4.1)
Here, $d$ is an integration constant. The energy density and pressure components can now be obtained by substituting $c_{13}=0$, $c_{14}=0$ and (4.1) nto the main equations (3)-(3), yielding
$\rho=\frac{1}{r^{7}(r^{4}-d^{2})}\Big\{(r^{4}-d^{2})[r^{4}(c_{2}+1)-d^{2}c_{2}]r\,b^{\prime}-c_{2}[(r^{8}-2d^{2}r^{4}-d^{4})b+2rd^{4}]\Big\}\,$,
$p_{t}=\frac{1}{2r^{3}(r^{4}\!-d^{2})}\bigg\{\!-2(r^{4}\!-d^{2})\big(c_{2}+\frac{1}{2}\big)r\,b^{\prime}+\left[(2c_{2}+1)r^{4}+2d^{2}\big(c_{2}-\frac{1}{2}\big)\right]b$
$-4d^{2}c_{2}r\bigg\},$
$p_{r}=\frac{1}{r^{7}(r^{4}-d^{2})}\Big\{(d^{2}-r^{4})c_{2}d^{2}r\,b^{\prime}+c_{2}\left[(r^{8}-2d^{2}r^{4}-d^{4})b+2rd^{4}\right]\Big\}\,.$ (4.2)
At this stage, by choosing an appropriate form for the wormhole shape function $b(r)$, it becomes straightforward to investigate the energy conditions for the wormhole solutions. In the following, we consider three types of $b(r)$: the power-law, logarithmic, and hyperbolic shape functions. [58].
4.1 Wormholes with power law shape function
As the first type, we choose the power-law shape function as [58]
$b(r)=r_{0}\left(\frac{r_{0}}{r}\right)^{n}\,$, (4.3)
We aim to identify regions in the parameter space of the solutions where the energy conditions are satisfied. Recall that the shape function $b(r)$ must also satisfy the conditions (3.2)-(3.4). It is straightforward to verify that, for the shape function (4.3) these conditions are satisfied when n>1n>-1. By substituting the shape function (4.3) into Eq. (4), the corresponding expressions for EC1, EC2, and EC3 are obtained as
$\!\!{\rm EC1}=\frac{1}{r^{7}(r^{4}\!-\!d^{2})}\Big\{2\left[c_{2}(n+1)+\frac{n}{2}\right]r_{0}^{n+1}d^{2}r^{4-n}-r_{0}^{n+1}\left[c_{2}(n+1)+n\right]r^{8-n}$
$~~~~~~~~\,-\left[r^{-n}(n-1)r_{0}^{n+1}+2r\right]c_{2}d^{4}\Big\}\,$,
$\!\!{\rm EC2}=\frac{1}{r^{7}(r^{4}\!-\!d^{2})}\Big\{3\Big[c_{2}\Big(n\!+\frac{5}{3}\Big)\!+\frac{1}{3}(n\!+\!1)\Big]r_{0}^{n+1}d^{2}r^{4-n}\!-[c_{2}(n\!+3)+n+1]r_{0}^{n+1}r^{8-n}$
$~~~~~~~~\,-2c_{2}\left[(n-1)d^{4}r^{-n}r_{0}^{n+1}-r^{9}+r^{5}d^{2}+2rd^{4}\right]\Big\}\,$,
$\!\!{\rm EC3}=\frac{1}{2r^{7}(r^{4}\!-\!d^{2})}\Big\{2d^{2}r_{0}^{n+1}r^{4-n}\Big[c_{2}(n+3)+\frac{1}{2}(n-1)\Big]-r_{0}^{n+1}r^{8-n}(n-1)$
$~~~~~~~~\,-2c_{2}d^{2}\left[d^{2}r_{0}^{n+1}r^{-n}(n-1)+2r^{5}+2rd^{2}\right]\Big\}$ (4.4)
At first glance, the wormhole solutions in this class depend on four parameters: $d$, $n$, $c_{2}$, and $r_{0}$. By setting $r_{0}=1$, we fix the unit of the radial coordinate $r$ to the size of the wormhole throat, leaving three parameters $d$, $n$, and $c_{2}$ in the solutions. In the following, we study the effect of these parameters on the energy conditions of the wormhole solutions.
In Fig. 1, the blue regions indicate areas in the parameter space of the wormhole solutions of the EA theory where ${\rm EC1}>0$, ${\rm EC2}>0$ and ${\rm EC3}>0$ simultaneously. In other words, for appropriate values of the parameters, the EA wormholes satisfy the WEC and, accordingly, the NEC. In particular, near the wormhole throat ($r_{0}=1$), the WEC and NEC are satisfied if $c_{2}>0$. Considering the constraints (2.7) on the coupling constants $c_{i}$ and noticing that $c_{13}=0$, one obtains $c_{2}>-\frac{2}{3}$, which is respected in our figures. However, our results indicate that the requirement of satisfying the energy conditions near the wormhole throat imposes a stronger constraint: $c_{2}>0$.
(a)(b)
Figure 1: The blue regions in these figures indicate where ${\rm EC1}>0$, ${\rm EC2}>0$, and ${\rm EC3}>0$ are satisfied simultaneously, meaning that the weak and null energy conditions are respected for the wormhole solutions in the EA theory. In these figures, we set $r_{0}=1$ and $d=5$. Fig. (a) shows that, by varying the coupling $c_{2}$, the WEC is satisfied near the wormhole throat if $c_{2}>0$ with $n=1$. In Fig. (b) the variation of the exponent $n$ versus $c_{2}$ is depicted for the radial distance $r=1.5$, and it is evident that the WEC is also satisfied for $c_{2}>0$.
In Fig. 2, by setting $c_{2}=\frac{1}{2}$, we show that the NEC (panel (a)) and WEC (panel (b)) are satisfied over the radial coordinate $r$ if the exponent $n$ in (4.3) lies within the interval indicated by the blue regions. A narrow gap appears around $r=2.5$, where neither the NEC nor WEC is satisfied. From Fig. 1, it is evident that in this narrow gap, the energy conditions are respected if $c_{2}$ takes negative values. Ignoring this small interval, one can conclude that the NEC and WEC are satisfied for the wormhole solutions of the EA theory when $c_{2}$ is positive. We also verified that variations in the parameter $d$ do not significantly affect these results.
(a)(b)
Figure 2: Satisfaction of the NEC and WEC for arbitrary values of the radial coordinate $r$ in the wormhole solutions of the EA theory is illustrated in these figures, plotted for $c_{2}=\frac{1}{2},d=5$, and $r_{0}=1$. In (a), the blue regions indicate where ${\rm EC2}>0$ and ${\rm EC3}>0$, showing that the NEC is satisfied for appropriate values of $n$ over nearly all radial distances outside the wormhole throat. A small gap appears around $r=2.5$ where the NEC is violated; however, within this gap, the NEC is satisfied for negative values of $c_{2}$ in the range $\frac{2}{3}<c_{2}<0$. A similar analysis is presented in (b) for the WEC, which is satisfied over a more limited region compared to the NEC.
4.2 Wormholes with logarithmic shape function
It is possible to consider the shape function in a logarithmic form [58]
$b(r)=r\,\frac{\ln r_{0}}{\ln r}\,.$ (4.5)
Noting the conditions (3.2)-(3.4), it is straightforward to show that the wormhole throat in this case must satisfy the bound $r_{0}>1$. By substituting the shape function (4.5) into (4), one can then obtain the corresponding expressions for the energy conditions as
$\!\!\!\!{\rm EC1}=\frac{1}{r^{6}(r^{4}\!-\!d^{2})(\ln r)^{2}}\bigg\{\left[(r^{8}-r^{4}d^{2}+2c_{2}d^{4})\ln r-(r^{4}-d^{2})\big(r^{4}(c_{2}+1)-d^{2}c_{2}\big)\right]\ln r_{0}$
$~~~~-2c_{2}(\ln r)^{2}d^{4}\bigg\},$
$\!\!\!\!{\rm EC2}=\frac{1}{r^{6}(r^{4}\!-\!d^{2})(\ln r)^{2}}\Big\{2c_{2}(\ln r)^{2}(r^{4}+d^{2})(r^{4}-2d^{2})$
$~~~~-\left[2c_{2}(r^{4}+d^{2})(r^{4}-2d^{2})\ln r+(r^{4}-d^{2})\big(r^{4}(c_{2}+1)-2d^{2}c_{2}\big)\right]\ln r_{0}\Big\},$
$\!\!\!\!{\rm EC3}=\frac{1}{2r^{6}(r^{4}\!-\!d^{2})(\ln r)^{2}}\bigg\{-4c_{2}(\ln r)^{2}d^{2}(r^{4}+d^{2})$
$~~~~+\left[\Big(2r^{8}\!+4\big(c_{2}-\frac{1}{2}\big)r^{4}d^{2}+4c_{2}d^{4}\Big)\!\ln r-2c_{2}d^{4}+2\big(c_{2}+\frac{1}{2}\big)r^{4}d^{2}-r^{8}\right]\!\ln r_{0}\bigg\}.$ (4.6)
To investigate the WEC and NEC, we examine whether there exist regions in the parameter space of these wormhole solutions where the above energy conditions are positive. Fig. 3 illustrates this parameter space. The blue regions in the figure indicate where EC1>>0, EC2>>0, and EC3>>0 simultaneously. It is evident from Fig. 3 that the NEC and WEC are violated for $c_{2}<0$. However, for suitable positive values within the bound $c_{2}>0$, both the NEC and WEC are satisfied at the wormhole throat and at large distances from it (except for a narrow interval). In other words, the satisfaction of the NEC and WEC for wormholes with the logarithmic shape function imposes a constraint on the coupling constant of the EA theory: $c_{2}>0$.
(a)(b)
Figure 3: Satisfaction of the energy conditions for arbitrary values of the radial coordinate $r$ in the case of wormhole solutions with a logarithmic shape function in the EA theory. The blue regions in (a) indicate where ${\rm EC2}>0$ and ${\rm EC3}>0$ simultaneously, implying that the NEC is satisfied. In (b), the blue regions correspond to the domain where all EC1, EC2, EC3 are positive, meaning that the WEC is respected. Our analysis shows that, for appropriate positive values of $c_{2}>0$, both the NEC and WEC are satisfied for nearly all radial distances outside the wormhole throat. These figures are plotted for $r_{0}=2$; however, we have verified that varying $r_{0}$ (within the bound $r_{0}>1$) does not significantly affect this overall behavior.
4.3 Wormholes with hyperbolic shape function
The last shape function considered in this section is the hyperbolic form. [58]
$b(r)=r_{0}\,\frac{\tanh r}{\tanh r_{0}}\,.$ (4.7)
In this case, all positive values of the wormhole throat radius $r_{0}>0$ satisfy the conditions (3.2)-(3.4). By substituting the above shape function into (4), we obtain the following expressions for the energy conditions:
$\!\!\!\!\!{\rm EC1}=\frac{1}{r^{7}\big(r^{4}-d^{2}\big)S_{0}C^{2}}\bigg\{r_{0}C_{0}\left[r(r^{4}-d^{2})\left(r^{4}(c_{2}+1)-c_{2}d^{2}\right)-SCc_{2}(r^{8}-2d^{2}r^{4}-d^{4})\right]$
$~~~\,-2S_{0}C^{2}c_{2}rd^{4}\bigg\},$
$\!\!\!\!{\rm EC2}=\frac{1}{r^{7}\big(r^{4}-d^{2}\big)S_{0}C^{2}}\Big\{r_{0}C_{0}\big[SC\left(-r^{8}(3c_{2}+1)+r^{4}d^{2}(5c_{2}+1)+2c_{2}d^{4}\right)$
$~~~~+r(r^{4}-d^{2})\big(r^{4}(c_{2}+1)-2c_{2}d^{2}\big)\big]+2rc_{2}S_{0}C^{2}(r^{4}+d^{2})(r^{4}-2d^{2})\Big\},$
$\!\!\!\!{\rm EC3}=\frac{1}{2r^{7}\big(r^{4}-d^{2}\big)S_{0}C^{2}}\Big\{r_{0}C_{0}\big[SC\left(r^{8}\!+r^{4}d^{2}(6c_{2}-1)+2c_{2}d^{4}\right)$
$~~~~+r(r^{4}\!-d^{2})\big(r^{4}\!-2c_{2}d^{2}\big)\big]-4rc_{2}d^{2}S_{0}C^{2}(r^{4}+d^{2})\Big\}\,$, (4.8)
where
$S=\sinh r,\quad S_{0}=\sinh r_{0},\quad C=\cosh r,\quad C_{0}=\cosh r_{0}\,.$ (4.9)
We performed a similar analysis and identified regions in the parameter space of these wormhole solutions where the energy conditions are satisfied. The results are presented in Fig. (4). As seen in this figure, when the EA coupling constant satisfies the bound $c_{2}>0$, the NEC is respected both near the wormhole throat and at large radial distances. In part (b) it is also evident that the WEC is satisfied around the wormhole throat for $c_{2}>0$.
Considering these results, we conclude that the satisfaction of the NEC and WEC for the wormhole solutions of the EA theory in the class $c_{13}=c_{14}=0$ imposes an additional constraint on the coupling constants, namely $c_{2}>0$. This bound can be combined with the constraints given in (2.7) (previously expressed as Eq. (2.7)) which were obtained from observational and theoretical considerations. Recall that for this class ($c_{13}=c_{14}=0$), Eq. (2.7) $c_{2}>-\frac{2}{3}$; thus, the wormhole condition $c_{2}>0$, imposes a stronger limitation on the value of $c_{2}$.
(a)(b)
Figure 4: The wormhole solutions with the hyperbolic shape function satisfy the NEC (panel a) and WEC (panel b) in the blue regions. By choosing a suitable value of the coupling constant of the EA theory such that $c_{2}>0$, the NEC is satisfied both at the wormhole throat and at large radial distances. The WEC is also respected around the wormhole throat for appropriate values within the bound $c_{2}>0$. In these figures we set $r_{0}=2$; however varying $r_{0}$ within the bound $r_{0}>0$ does not alter this overall behavior.
5 Class II: $c_{2}=0$, and $c_{14}=0$
Another approach to solving the main equations (3)-(3) is to set the coupling constants of the EA theory to $c_{2}=0$ and $c_{14}=0$. In this case, the aether field is expressed by the function [56]
$a(r)=\frac{e}{r^{2}}\,$, (5.1)
where $e$ is an integration constant. By substituting $c_{2}=0$ and $c_{14}=0$ together with (5.1) into the field equations (3)-(3), one obtains the expressions for the energy density $\rho$, and the radial and transverse pressure components $p_{r}$ and $p_{t}$, as
$\rho$ $=$ $\frac{1}{r^{11}(r^{4}\!-\!e^{2})}\Big\{r(r^{4}\!-\!e^{2})\left[r^{8}+e^{2}(c_{3}-c_{4})(r^{4}-e^{2})\right]b^{\prime}$
$+$ $(c_{3}-c_{4})\left[(-r^{12}-3e^{2}r^{8}+7e^{4}r^{4}-e^{6})b+r^{5}(r^{8}+2e^{2}r^{4}-5e^{4})\right]\Big\}\,$,
$p_{r}$ $=$ $\frac{1}{r^{11}(r^{4}\!-\!e^{2})}\Big\{-(c_{3}-c_{4})re^{4}(r^{4}-e^{2})b^{\prime}+(c_{3}-c_{4})r^{5}(r^{8}-4e^{2}r^{4}+e^{4})$
$+$ $\left[-(c_{3}-c_{4}+1)r^{12}+4e^{2}\big(c_{3}-c_{4}+\frac{1}{4}\big)r^{8}-e^{6}(c_{3}-c_{4})\right]b\Big\}\,$,
$p_{t}$ $=$ $\frac{1}{2r^{7}(r^{4}\!-\!e^{2})}\Big\{-r(r^{4}-e^{2})\big[(c_{3}-c_{4}+1)r^{4}-(c_{3}-c_{4})e^{2}\big]b^{\prime}$ (5.2)
$+$ $8(c_{3}-c_{4})\big(r^{4}-\frac{3}{2}e^{2}\big)e^{2}r+\left[(c_{3}-c_{4}+1)r^{8}-10e^{2}\big(c_{3}-c_{4}+\frac{1}{10}\big)r^{4}\right$.
$+$ $\left.13e^{4}(c_{3}-c_{4})\right]b\Big\}\,.$
At this stage, we substitute three types of shape functions, power-law, logarithmic, and hyperbolic, into the above equations and investigate the corresponding energy conditions for the wormhole solutions of the EA theory in this case.
5.1 The power law shape function
We begin our analysis in this section with the power-law shape function $b(r)=r_{0}(r_{0}/r)^{n}$. By substituting this function into the energy density and pressure expressions (5), the corresponding energy conditions can be obtained as follows:
$\!\!\!\!\!\!\!{\rm EC1}=\frac{1}{r^{11}(r^{4}\!-\!e^{2})}\Big\{-e^{2}[(c_{3}-c_{4})(n+3)-n]r_{0}^{n+1}r^{8-n}-r_{0}^{n+1}r^{12-n}[n+(c_{3}-c_{4})]$
$~~~~~-(c_{3}\!-\!c_{4})\left[-2e^{4}\big(n+\frac{7}{2}\big)r_{0}^{n+1}r^{4-n}+e^{6}r_{0}^{n+1}r^{-n}(n+1)r^{13}-2r^{9}e^{2}+5r^{5}e^{4}\right]\Big\}\,$,
$\!\!\!\!\!\!\!{\rm EC2}=\frac{1}{r^{11}(r^{4}\!-\!e^{2})}\Big\{\!-e^{2}[(c_{3}-c_{4})(n\!-\!1)-n\!-\!1]r_{0}^{n+1}r^{8-n}\!-r_{0}^{n+1}r^{12-n}[n\!+\!1\!+\!2(c_{3}\!-c_{4})]$
$~~~~~-2(c_{3}\!-\!c_{4})\left[-\frac{3}{2}e^{4}r_{0}^{n+1}r^{4-n}\big(n+\frac{7}{3}\big)+e^{6}r_{0}^{n+1}r^{-n}(n+1)-r^{13}+r^{9}e^{2}+2r^{5}e^{4}\right]\Big\}\,$,
$\!\!\!\!\!\!\!{\rm EC3}=\frac{1}{2r^{11}(r^{4}\!-\!e^{2})}\Big\{(c_{3}\!-\!c_{4}\!-1)(n\!-\!1)r_{0}^{n+1}r^{12-n}\!-[4(n\!+4)(c_{3}\!-\!c_{4})-n+1]e^{2}r_{0}^{n+1}r^{8-n}$
$~~~~~-2(c_{3}\!-\!c_{4})\left[-\frac{5}{2}\big(n\!+\!\frac{27}{5}\big)e^{4}r_{0}^{n+1}r^{4-n}\!+e^{6}(n\!+\!1)r_{0}^{n+1}r^{-n}\!-r^{13}\!-6r^{9}e^{2}\!+11r^{5}e^{4}\right]\!\Big\}.$
Note that by setting $r_{0}=1$, only two parameters $e$ and $n$ and the combination $c_{3}-c_{4}$ remain for the solutions in this class. We are now in a position to investigate the energy conditions and identify regions in the parameter space where the NEC and WEC are satisfied. Due to the constraints (3.2)-(3.4) the parameter $n$ in the power-law shape function (4.3) should satisfy the bound n>1n>-1, which has been taken into account in our analysis.
Figure 5: The inequalities ${\rm EC1}>0$, ${\rm EC2}>0$, and ${\rm EC3}>0$ are satisfied simultaneously in the blue regions, indicating that both the WEC and NEC are respected for the wormhole solutions with the power-law shape function in the EA theory. Here, we set $r_{0}=1$, $n=1$, and $e=3$. It is evident that, except for a narrow interval around $r=2$ (twice the throat radius), the WEC and NEC are satisfied throughout the spacetime for the wormhole solutions of the EA theory, provided that the combination $c_{3}-c_{4}>0$ .
The blue region in Fig. 5 indicates where ${\rm EC1}>0$, ${\rm EC2}>0$ and ${\rm EC3}>0$ simultaneously, meaning that the WEC and NEC are satisfied. From Fig. 5 it is evident that if the coupling constants $c_{3}$ and $c_{4}$ satisfy the bound $c_{3}-c_{4}>0$, the WEC is respected around the wormhole throat (located at $r_{0}=1$). Moreover, except for a narrow interval around $r=2$, the WEC and NEC are satisfied for all remaining $r>r_{0}$, when $c_{3}-c_{4}$ is positive.
To explore the effect of the exponent $n$ on the WEC and NEC, we plot Fig. 6. It is clear from (a) that as $n$ decreases, the required value of $c_{3}-c_{4}$ increases. However, for any $n$ satisfying n>1n>-1, the WEC and NEC are preserved if $c_{3}-c_{4}>0.5$. As shown in part (b) of Fig. 6, by fixing $c_{3}-c_{4}=1$, the WEC and NEC are satisfied for almost all radial distances, except for a narrow range around $r=2$ for any exponent n>1n>-1.
We also performed a similar analysis for the satisfaction of $EC2>0$ and $EC>03$ (corresponding to the NEC only, where the WEC may be violated). The results show no significant differences. In other words, the wormhole solutions of the EA theory with $c_{2}=0$ and $c_{14}=0$ satisfy both the WEC and NEC if the combination $c_{3}-c_{4}$ satisfies the bound $c_{3}-c_{4}>\frac{1}{2}$. We also verified that varying the parameter $e$ does not significantly affect these results.
(a)(b)
Figure 6: The effect of $n$ on the satisfaction of the NEC and WEC in the wormhole solutions of the EA theory is illustrated in these figures. The plots are generated by setting $r_{0}=1,e=5$. Part (a) plotted for $r=3$ shows that for large values of $n$, the WEC and NEC are satisfied when $c_{3}-c_{4}>0$; as $n$ decreases, the required value of $c_{3}-c_{4}$ must increase. However, if $c_{3}-c_{4}>0.5$ the WEC and NEC are satisfied for all allowed values of $n$. In part (b) we sett $c_{3}-c_{4}=1$ and it is evident that the WEC and NEC are satisfied for almost all radial distances, independently of the value of $n$.
5.2 The logarithmic shape function
The energy conditions for the wormhole solutions in this case can be obtained by substituting the logarithmic shape function (4.5) into the energy density and pressure expressions (5). The result reads
$\!\!\!\!\!\!{\rm EC1}=\frac{1}{r^{10}(r^{4}\!-e^{2})(\ln r)^{2}}\biggl\{\!\Big\{\!\!-\!r^{4}\Big[(c_{3}-c_{4}-1)r^{8}\!+2e^{2}\Big(c_{3}\!-\!c_{4}+\frac{1}{2}\Big)r^{4}\!-5e^{4}(c_{3}\!-\!c_{4})\Big]\ln r$
$~~~-r^{12}\!-e^{2}(c_{3}-c_{4}-1)r^{8}\!+2e^{4}(c_{3}-c_{4})r^{4}\!-e^{6}(c_{3}-c_{4})\Big\}\ln r_{0}$
$~~~+r^{4}(c_{3}-c_{4})\big(r^{8}\!+2e^{2}r^{4}-5e^{4}\big)(\ln r)^{2}\biggr\}\,$,
$\!\!\!\!\!\!{\rm EC2}=\frac{1}{r^{10}(r^{4}\!-e^{2})(\ln r)^{2}}\Big\{\!\big[\!-2r^{4}(c_{3}\!-\!c_{4})(r^{4}+e^{2})(r^{4}-2e^{2})\ln r-r^{12}\!-e^{2}(c_{3}\!-\!c_{4}-1)r^{8}$
$~~~+3e^{4}(c_{3}-c_{4})r^{4}\!-2e^{6}(c_{3}-c_{4})\big]\ln r_{0}+2r^{4}(c_{3}-c_{4})(r^{4}+e^{2})(r^{4}-2e^{2})(\ln r)^{2}\Big\}\,$,
$\!\!\!\!\!\!{\rm EC3}=\frac{1}{2r^{10}(r^{4}\!-e^{2})(\ln r)^{2}}\biggl\{\!\Big\{\!\!-2r^{4}\Big[(c_{3}-c_{4}-1)r^{8}\!+6e^{2}\Big(c_{3}-c_{4}+\frac{1}{6}\Big)r^{4}\!-11e^{4}(c_{3}\!-\!c_{4})\Big]$
$~~~\times\ln r+(r^{4}-e^{2})\big[(c_{3}-c_{4}-1)r^{8}\!-3e^{2}(c_{3}-c_{4})r^{4}\!+2e^{4}(c_{3}-c_{4})\big]\Big\}\ln r_{0}$
$~~~+2r^{4}(c_{3}-c_{4})\big(r^{8}\!+6e^{2}r^{4}-11e^{4}\big)(\ln r)^{2}\biggr\}\,.$ (5.4)
Here, the bound $r_{0}>1$ must be satisfied at the wormhole throat due to the conditions (3.2)-(3.4). Similar to the previous cases, we focus on solutions that satisfy the NEC and WEC. In Fig. 7 we depict the region in which all energy conditions in (5.2) are positive, indicating that the WEC is satisfied. It is evident that, for wormhole solutions with a logarithmic shape function, the WEC is satisfied at the throat and almost everywhere, provided appropriate values for the couplings within the bound $c_{3}-c_{4}>0$ are chosen.
We also examine the satisfaction of EC2>>0 together with EC3>>0 (corresponding to the NEC), , and the results are the same as in Fig. 7. In other words, the satisfaction of the NEC and WEC for wormholes with the logarithmic shape function imposes an additional constraint on the values of the coupling constants in EA theory: $c_{3}-c_{4}>0$.
Figure 7: Wormhole solutions of the EA theory with a logarithmic shape function satisfy the WEC and NEC in the blue region. By choosing a suitable value for the coupling within the bound $c_{3}-c_{4}>0$, the WEC and NEC are satisfied at the wormhole throat (located at $r_{0}=2$) as well as at large distances. This figure is depicted for $e=5$; however, we verified that other values of $e$ and $r_{0}$ do not alter this overall behavior.
5.3 The hyperbolic shape function
Considering the hyperbolic shape function for the wormhole solutions in this case, the energy conditions can be directly obtained by substituting (4.7) into (5). The result is
$\!\!\!\!{\rm EC1}=\frac{1}{r^{11}(r^{4}-e^{2})\tanh r_{0}}\bigg\{\!-r_{0}r(r^{4}-e^{2})[r^{8}+(c_{3}-c_{4})e^{2}r^{4}\!-(c_{3}-c_{4})e^{4}](\tanh r)^{2}$
$~~~~-r_{0}(c_{3}-c_{4})(r^{12}+3e^{2}r^{8}-7e^{4}r^{4}+e^{6})\tanh r+r\Big\{r^{4}(c_{3}-c_{4})(r^{8}\!+2e^{2}r^{4}\!-5e^{4})$
$~~~~\times\tanh r_{0}+r_{0}(r^{4}-e^{2})\left[r^{8}+(c_{3}-c_{4})e^{2}r^{4}-(c_{3}-c_{4})e^{4}\right]\Big\}\bigg\},$
$\!\!\!\!{\rm EC2}=\frac{1}{r^{11}(r^{4}\!-e^{2})S_{0}C^{2}}\bigg\{\!\Big\{\!-2\Big[\Big(c_{3}-c_{4}+\frac{1}{2}\Big)r^{12}\!-\frac{1}{2}e^{2}(c_{3}-c_{4}+1)r^{8}\!-\frac{7}{2}e^{4}(c_{3}\!-\!c_{4})r^{4}$
$~~~~+e^{6}(c_{3}-c_{4})\Big]SC+r(r^{4}-e^{2})[r^{8}+e^{2}(c_{3}\!-\!c_{4})r^{4}\!-2e^{4}(c_{3}\!-\!c_{4})]\Big\}r_{0}C_{0}$
$~~~~+2r^{5}(c_{3}-c_{4})C^{2}S_{0}(r^{4}+e^{2})(r^{4}-2e^{2})\bigg\},$
$\!\!\!\!{\rm EC3}=\frac{1}{2r^{11}(r^{4}\!-e^{2})S_{0}C^{2}}\Big\{\!-\!\big\{\!\big[(c_{3}\!-\!c_{4}\!-1)r^{12}\!+(16(c_{3}-c_{4})+1)e^{2}r^{8}\!-27e^{4}(c_{3}\!-\!c_{4})r^{4}$
$~~~~+2e^{6}(c_{3}-c_{4})\big]SC+r(r^{4}\!-e^{2})[(c_{3}-c_{4}-1)r^{8}\!-3(c_{3}\!-\!c_{4})e^{2}r^{4}\!+2e^{4}(c_{3}\!-\!c_{4})]\big\}r_{0}C_{0}$
$~~~~+2r^{5}(c_{3}-c_{4})C^{2}S_{0}(r^{8}+6e^{2}r^{4}-11e^{4})\Big\}\,$, (5.5)
where the definitions have been used for the sake of brevity. Remember that in this case $r_{0}>0$ must be satisfied due to the conditions (3.2)-(3.4). The region in which the WEC is respected is depicted in Fig. 8. It is evident that by selecting appropriate values within the bound $c_{3}-c_{4}>1$, the wormhole solutions satisfy the WEC at the wormhole throat and at large distances, except in a narrow range near the throat. The same behavior is observed for the NEC for the wormhole solutions in this case.
Figure 8: The WEC and NEC are respected in the blue regions for the wormhole solutions of the EA theory with hyperbolic shape function by choosing appropriate values for the couplings in the bound $c_{3}-c_{4}>0$. In this figure we set $r_{0}=2$ and $e=5$, however we checked varying $e$ and $r_{0}$ does not change this total behavior.
We now summarize the results of this section. We found that the EA theory, for $c_{2}=0$ and $c_{14}=0$, admits wormhole solutions with power-law, logarithmic, and hyperbolic shape functions. The satisfaction of the NEC and WEC for these wormholes imposes additional constraints on the values of the couplings as follows:
  • $c_{3}-c_{4}>\frac{1}{2}$ for wormhole solutions with the power-law shape function,
  • $c_{3}-c_{4}>0$ for wormhole solutions with logarithmic and hyperbolic shape function.
By merging these two constraints, we deduce that for all wormhole solutions in this class, the NEC and WEC are satisfied if $c_{3}-c_{4}>\frac{1}{2}$. Note that in (2.7) there is no constraint on the combination of $c_{3}$ and $c_{4}$, so the wormhole constraint $c_{3}-c_{4}>\frac{1}{2}$ obtained here puts a stronger limitation on the values of these couplings.
6 Class III: $c_{2}=-c_{13}\neq 0$, and $c_{14}=0$
It is also possible to solve the main equations (3)-(3) by setting the coupling constants as $c_{2}=-c_{13}\neq 0$ and $c_{14}=0$. with this choice the aether field can be obtained as [56]
$a(r)=-\frac{1}{c_{2}r}\sqrt{c_{2}r(jc_{2}-r)}\,$, (6.1)
where $j$ is an integration constant. By substituting the above function and the values $c_{2}=-c_{13}$ and $c_{14}=0$ into the main equations, we obtain the corresponding energy density and pressure components as
$\rho$ $=$ $\frac{1}{c_{2}r^{5}}\Big\{r\big[(c_{2}^{2}+3c_{2}+1)r^{2}-2jc_{2}(c_{2}+1)r+j^{2}c_{2}^{2}\big]b^{\prime}$
$+$ $\big[-(c_{2}+1)r^{2}+2jc_{2}(c_{2}+1)r-j^{2}c_{2}^{2}\big]b-c_{2}r^{3}(c_{2}+1)\Big\}\,$,
$p_{r}$ $=$ $\frac{1}{c_{2}r^{5}}\Big\{r(r-c_{2}j)[(c_{2}+1)r-c_{2}j]b^{\prime}+c_{2}r^{3}(c_{2}+1)$
$+$ $[-r^{2}(c_{2}^{2}+3c_{2}+1)+rc_{2}j(c_{2}+2)-j^{2}c_{2}^{2}]b\Big\}\,$,
$p_{t}$ $=$ $-\frac{1}{2r^{4}}\bigg\{\left[(c_{2}+2)r-\frac{1}{2}jc_{2}\right](b^{\prime}r-b)\bigg\}\,.$ (6.2)
Similar to the previous sections, we now consider three different shape functions for the wormhole solutions of the EA theory in this case and examine the energy conditions for them. Recalling the constraint (2.7) among the coupling constants, it is straightforward to see that the bound $c_{2}>-1$ must be satisfied for the solutions in this class, and this restriction has been taken into account in the analysis and figures presented in this section.
6.1 The power law shape function
In the case of wormhole solutions with the power-law shape function (4.3), and taking into account the energy density and pressure components given in (6), one can express the energy conditions as follows
$\!\!{\rm EC1}=\frac{1}{c_{2}r^{5}}\Big\{-[nc_{2}^{2}+(3n+1)c_{2}+n+1]r^{2-n}r_{0}^{n+1}$
$~~~~~-c_{2}[-2j(c_{2}+1)(n+1)r^{1-n}r_{0}^{n+1}+j^{2}c_{2}(n+1)r^{-n}r_{0}^{n+1}+(c_{2}+1)r^{3}]\Big\}\,$,
$\!\!{\rm EC2}=-\frac{1}{c_{2}r^{5}}\bigg\{2(n+1)\left[-\frac{3}{2}c_{2}j\left(c_{2}+\frac{4}{3}\right)r^{1-n}+\left(\frac{1}{2}c_{2}^{2}+2c_{2}+1\right)r^{2-n}+c_{2}^{2}j^{2}r^{-n}\right]$
$~~~~~\times r_{0}^{n+1}\bigg\},$
$\!\!{\rm EC3}=\frac{1}{4c_{2}r^{5}}\bigg\{-2\left[(n-1)c_{2}^{2}+4c_{2}n+2(n+1)\right]r_{0}^{n+1}r^{2-n}$
$~~~~~-4c_{2}\!\left[-\frac{7}{4}j(n\!+\!1)\left(c_{2}\!+\frac{8}{7}\right)r_{0}^{n+1}r^{1-n}\!+c_{2}j^{2}(n\!+\!1)r_{0}^{n+1}r^{-n}\!+r^{3}(c_{2}\!+\!1)\right]\!\bigg\}.$
From the blue-colored region in Fig. 9, it is evident that the NEC and the WEC are satisfied for the wormhole solutions in this case when the coupling constant $c_{2}$ takes negative values. In this figure it is clearly seen that both the NEC and WEC are satisfied around the wormhole throat (which is located at $r_{0}=1$) when $1<c_{2}<0$. As an important feature of this class, we observe that the energy conditions are also satisfied throughout the entire spacetime, when the coupling constant $c_{2}$ lies within the range $1<c_{2}<0$.
Figure 9: The blue region in this figure shows the domain where the NEC and WEC are satisfied. One can see that the energy conditions are respected throughout the entire spacetime for the wormholes with a power-law shape function in this class, provided that the coupling constant $c_{2}$takes appropriate negative values. In this figure, we set $r_{0}=1,n=-3/4$ and $j=3$.
The role of exponent $n$ in satisfying the NEC and WEC, is explored in Fig. 10 where the parameters $c_{2}$ and $r$ are varied with respect to $n$ to illustrate the regions in which the energy conditions are satisfied. It can be seen from Fig. 10(a) that the NEC is satisfied for a wider range of exponent values in the bound 1<n-1<n, provided that the coupling constant $c_{2}$ lies within the range $1<c_{2}<0$. This behavior can be understood more clearly from Fig. 10(b), where the NEC region is plotted by fixing the coupling constant at $c_{2}=-1/2$ and varying $n$ with respect to $r$. It is evident from this figure that the NEC is satisfied throughout the entire spacetime when $n$ takes small positive values within the bound 1<n-1<n. In Figs. 10(c) and 10(d) we perform a similar analysis for the WEC. Our results are qualitatively the same as for the NEC, except for minor differences in the size of the corresponding regions.
(a)(b)(c)(d)
Figure 10: The effect of the exponent $n$ on the energy conditions is illustrated in this figure. Here, we set $r_{0}=1$ and $j=3$. (a) shows the region in which the NEC is satisfied for $r=3$. In (b) by setting $c_{2}=-1/2$, we demonstrate the satisfaction of the NEC throughout the radial coordinate $r$. In (c) and (d) we perform a similar analysis for the WEC.
6.2 The logarithmic shape function
To find the energy conditions for wormhole solutions of the EA theory with logarithmic shape function, it is sufficient to substitute (4.5) into the energy density and pressures given by (6). Doing this, one finds
$\!\!\!\!{\rm EC1}=\frac{1}{c_{2}(\ln r)^{2}r^{4}}\bigg\{\left[c_{2}(c_{2}+2)r^{2}\ln r-c_{2}^{2}(r-j)^{2}+c_{2}(2jr-3r^{2})-r^{2}\right]\ln r_{0}$
$~~~~-c_{2}(c_{2}+1)(\ln r)^{2}r^{2}\bigg\},$
$\!\!\!\!{\rm EC2}=-\frac{1}{c_{2}(\ln r)^{2}r^{4}}\bigg\{\left[c_{2}^{2}(r-j)(r-2j)+4c_{2}r(r-j)+2r^{2}\right]\ln r_{0}\bigg\}\,$,
$\!\!\!\!{\rm EC3}=\frac{1}{4c_{2}(\ln r)^{2}r^{4}}\Big\{\Big[4c_{2}(c_{2}+2)r^{2}\ln r+c_{2}^{2}(-4j^{2}\!+7jr-2r^{2})-8c_{2}r(r\!-j)-4r^{2}\Big]\ln r_{0}$
$~~~~-4c_{2}(c_{2}+1)r^{2}(\ln r)^{2}\Big\}\,.$ (6.3)
Fig. 11 (a) and (b) show the satisfaction of NEC and WEC respectively. One can easily observe that if the coupling $c_{2}$ lies in the range $1<c_{2}<0$, the NEC and WEC are respected at the wormhole throat located at $r_{0}=2$. The energy conditions could also be respected throughout the whole space, by choosing small negative values for $c_{2}$.
(a)(b)
Figure 11: In the blue region in (a) the NEC is respected by the wormholes of the EA theory with logarithmic shape function. (b) shows the same region for the WEC. Here, we set $r_{0}=2$ and $j=1$, however this behavior does not affected totally by varying $j$ and $r_{0}$.
6.3 The hyperbolic shape function
For the wormhole solutions with the hyperbolic shape function in this class, the energy conditions can be obtained by inserting the hyperbolic shape function (4.7) into the energy density and pressures (6),
$\!\!\!{\rm EC1}=\frac{1}{c_{2}r^{5}\tanh r_{0}}\biggl\{-r_{0}r\left[(c_{2}^{2}+3c_{2}+1)r^{2}-2jc_{2}(c_{2}+1)r+c_{2}^{2}j^{2}\right](\tanh r)^{2}$ (6.4)
$~~~~-r_{0}\left[(c_{2}+1)r^{2}-2c_{2}(c_{2}+1)jr+c_{2}^{2}j^{2}\right]\tanh r$
$~~~~+r\Big\{-c_{2}(c_{2}+1)r^{2}\tanh r_{0}+r_{0}\big[(c_{2}^{2}+3c_{2}+1)r^{2}-2c_{2}(c_{2}+1)jr+c_{2}^{2}j^{2}\big]\Big\}\biggr\}\,$,
$\!\!\!{\rm EC2}=\frac{1}{c_{2}r^{5}C^{2}S_{0}}\bigg\{\left[(c_{2}^{2}+4c_{2}+2)r^{2}-3c_{2}j\Big(c_{2}+\frac{4}{3}\Big)r+2c_{2}^{2}j^{2}\right](r-SC)C_{0}r_{0}\bigg\}\,$,
$\!\!\!{\rm EC3}=\frac{1}{4c_{2}r^{5}C^{2}S_{0}}\bigg\{2\Big\{\Big[(c_{2}^{2}-2)r^{2}+\frac{7}{2}c_{2}j\Big(c_{2}+\frac{8}{7}\Big)r-2c_{2}^{2}j^{2}\Big]SC$
$~~~~+r\Big[(c_{2}^{2}+4c_{2}+2)r^{2}-\frac{7}{2}c_{2}j\Big(c_{2}+\frac{8}{7}\Big)r+2c_{2}^{2}j^{2}\Big]\Big\}r_{0}C_{0}-4c_{2}(c_{2}+1)r^{3}S_{0}C^{2}\bigg\}.$
The same analysis shows that, these wormhole solutions respect the NEC and WEC on the throat $r_{0}$ and also throughout the whole space outside the $r_{0}$, if the coupling constant of the EA theory chosen adequately in the range $1<c_{2}<0$. Fig. 12 shows the parameter space of these wormhole solutions. Similar to the previous cases, in order to satisfy the NEC and WEC faraway from the throat, the coupling $c_{2}$ should take small negative values.
(a)(b)
Figure 12: Satisfaction of the energy conditions in the case of wormholes with hyperbolic shape function is depicted. In the blue region of (a) the NEC is respected. (b) shows the same region for the WEC. Here, we set $r_{0}=2$ and $j=1$.
We can summarized our result in this section as follows. The EA theory with the values of couplings $c_{2}=-c_{13}\neq 0$ and $c_{14}=0$, admits wormhole solutions with power law, logarithmic and hyperbolic shape functions. We found a distinct feature for the EA theory in class III that is by choosing appropriate parameters, the wormholes solutions respect the energy conditions throughout the whole space. Satisfaction of the NEC and WEC for wormholes in this class put an additional constraint on the coupling constant as $c_{2}<0$. This bound may be added to the constraints (2.7), which for this class leads to the bound $1<c_{2}$. Combining our wormhole bound with this bound, one concludes that the value of coupling should by in the range $1<c_{2}<0$ .
Before concluding, it is worth to mention that there are some other cases in which the main equations (3)-(3) can be solved. For instance, we find wormhole solutions for the classes: IV) $c_{2}\neq 0$, $c_{14}=0$, and $c_{13}\neq 0$ , V) $c_{2}=0$, $c_{14}=0$, and $c_{13}\neq 0$, and so on. We also checked that the energy conditions for the wormholes in these classes could be respected if we choose the corresponding parameters adequately. However, for these classes, there are numerous parameters in the parameter space of the wormhole solutions, so that we could not find any meaningful constraint between them by investigating the energy conditions.
7 Summary
In the context of Einstein gravity, wormhole geometries typically violate the energy conditions due to the flare–out condition at the wormhole throat [59]. In order to construct wormhole geometries that can be supported by ordinary (non-exotic) matter, we considered traversable wormholes within the framework of EA theory. By adopting an anisotropic energy–momentum tensor, we solved the equations of motion of the EA theory for specific combinations of the coupling constants. We then derived the corresponding energy density and pressure components for the wormhole solutions belonging to the three identified classes in the EA theory, as follows:
  • Class I:     $c_{2}\neq 0$, $c_{13}=0$, and $c_{14}=0\,$,
  • Class II:    $c_{2}=0$, and $c_{14}=0\,$,
  • Class III:   $c_{2}=-c_{13}\neq 0$, and $c_{14}=0\,$.
From previous observational and theoretical studies, several constraints on the coupling constants $c_{i}$ of the EA theory have been obtained, as follows [53, 54, 55]
$0\leq c_{14}<2\,,\qquad 2+c_{13}+3c_{2}>0\,,\qquad 0\leq c_{13}<1\,.$ (7.1)
In this work, we investigate the NEC and WEC for wormhole geometries by considering three different types of shape functions, as follows
  • Power law shape function:     $b(r)=r_{0}\left(r_{0}/r\right)^{n}\,$,
  • Logarithmic shape function:     $b(r)=r\left(\ln r_{0}/\ln r\right)\,$,
  • Hyperbolic shape function:     $b(r)=r_{0}\left(\tanh r/\tanh r_{0}\right)\,$,
for each classes of coupling values. For these nine cases, we found that the wormhole solutions in EA theory satisfy the NEC and WEC at the wormhole throat and at large distances, provided that the parameters $r_{0}$, $n$, and $c_{i}$ are appropriately chosen. In the case of wormhole geometries belonging to Class III, we observed that the energy conditions can be satisfied not only at the wormhole throat but also throughout the entire spacetime.
Our results show that satisfying the NEC and WEC for the wormhole solutions can impose additional constraints on the values of the coupling constants in EA theory, as follows:
In Class I, we found that the coupling constant $c_{2}$ must satisfy the bound $c_{2}>0$. This constraint is more restrictive than the previous bound $c_{2}>-\frac{2}{3}$ obtained from (7.1) for this class.
In Class II, satisfying the NEC and WEC imposes a constraint on the combination of coupling constants $c_{3}$ and $c_{4}$ as: $c_{3}-c_{4}>\frac{1}{2}$. This represents a new constraint compared to (7.1).
Finally, in Class III, we found that the wormhole solutions satisfy the energy conditions if the coupling constant $c_{2}$ lies in the range $c_{2}<0$. This wormhole bound can be combined with the previous bound $1<c_{2}$, obtained from (7.1) for this class, to give a single constraint $1<c_{2}<0$, which places a stronger limitation on the value of the coupling $c_{2}$, in the EA theory.